Can you give me the solution to this problem please?
0.04(y-2)+0.02y=0.16y-0.5

Answer by shweta(56)

0.04(y-2)+0.02y=0.16y-0.5
0.04y-0.08 + 0.02y = 0.16y- 0.5
0.06y - 0.16y = 0.08 -0.5
-0.10y = -0.42
y =0.42/0.10
y=42/10
y=4.2